# A Survey of Graph Unlearning

### v3

@@ 3.2.1 Convex Setting
The paper [41] introduces a graph embedding technique called Graph Scattering Transform (GST). GST relies on a collection of multiresolution graph wavelets, a pointwise nonlinear activation function, and a low-pass operator. Building upon GST, [20] propose an interesting graph unlearning approach, aptly named Unlearning GST. By utilizing GST embeddings, Unlearning GST defines an unlearning model denoted as $\mathcal{M}^{\prime}$, which updates the trained model from $W$ to $W^{\prime}$. The latter represents an approximation of the unique optimizer of $(L(W,\mathcal{D}^{\prime}))$. The derivation of $W^{\prime}$ is as follows:
To begin, $H_{W}$ is defined as the Hessian of $(L(.,\mathcal{D}^{\prime}))$ at $W$ and is computed as:
@@
Using the obtained gradient difference, $W^{\prime}$ is derived as follows:
@@
Thus, $W^{\prime}$ represents the updated model obtained through this unlearning process. However, the proposed approach has a limitation related to the computation of the wavelet transform, which is cubic for node removal and quadratic for node feature removal.
@@
The authors in [33] introduce another unlearning method, Graph Influence Function (GIF). GIF aims to model the influence of each training data point on the model with respect to various performance criteria and subsequently eliminates the negative impact. The method incorporates an additional loss term that considers the influence of neighbors and estimates parameter changes in response to a $\epsilon-$mass perturbation in deleted data. Another similar approach is presented by Cong et al. in their work [34], where they introduce “Projector”. This method aims to remove the impact of deleted data from the learned model by projecting the weight parameters to a different subspace that is unrelated to the deleted data. Specifically, Projector is designed to work with linear GNNs, and it demonstrates that all gradients lie within the linear span of all node features. Leveraging this insight, it performs unlearning using an orthogonal projection represented as a weighted combination of the remaining node features. By applying this technique, the learned model can effectively adapt to the removal request and update itself while minimizing the impact of the deleted data.
@@ 3.2.2 Non-Convex Setting
In non-convex setting, a notable contribution in graph-level unlearning context can found in [42]. This approach offers a method for certified removal of deleted data from a trained model. Operating within the non-convex setting, the method applies the Newton update, also known as the influence function, to the gradient influence of the removed data point. The paper goes on to establish robust theoretical bounds on the security parameters, providing rigorous proof of the certified removal’s effectiveness.
@@
In the pursuit of approximate unlearning within the non-convex setting, EraEdge is introduced in [31]. EraEdge specifically targets edge unlearning in graphs. To achieve this, EraEdge focuses on reversing the influence of the edge that needs to be removed from the new model $W^{\prime}$, leveraging the concept of influence functions [43]. The authors propose an estimation of the influence function by upweighting the set of all affected nodes and subsequently calculate the reverse of the Hessian matrix multiplied by the gradient vector to obtain the influence. To address computational complexity, they employ conjugate gradient optimization, effectively reducing the computational burden.
@@
The unlearning process, as explored in [44], can potentially harm the performance of the underlying predictive model. In an effort to mitigate this issue and avoid updating the trained model, [21] introduces GNNDELETE. This approach formalizes two essential properties for the GNN deletion method: the deleted edge consistency and neighborhood influence. The deleted edge consistency ensures that the predicted probabilities for deleted edges in the unlearned model stay similar to those for nonexistent edges. Similarly, the neighborhood influence guarantees that predictions in the local vicinity of the deletion retain their original performance and remain unaffected by the removal. In pursuit of efficiency and scalability, GNNDELETE employs a layer-wise deletion operator to modify a pre-trained GNN model. Upon receiving deletion requests, GNNDELETE freezes the existing model weights and introduces small, shared weight matrices across the nodes in the graph. GNNDELETE guarantees strong performance by ensuring that the difference between node representations obtained from the trained model $\mathcal{M}$ and those revised by GNNDELETE $\mathcal{M}^{\prime}$ remains theoretically bounded.
In the knowledge graph setting, the authors in [35] introduce FedLU, a federated learning framework for knowledge graphs (KG). The knowledge graph setting involves complex and diverse data, and FedLU addresses the challenges arising from data heterogeneity. One of the key contributions of FedLU is the introduction of a mutual knowledge distillation method. This method plays a crucial role in coping with the drift that can occur between local optimization and global convergence due to the heterogeneity present in the KG data. By utilizing mutual knowledge distillation, FedLU aims to harmonize the information learned across different local models and ensure more consistent global convergence. This not only helps in maintaining model performance but also enables better collaboration and sharing of knowledge across the participating entities in the federated learning setting. Another significant aspect of FedLU is the proposal of a KG embedding unlearning method. Knowledge forgetting is an essential requirement in dynamic knowledge graphs where information needs to be updated or removed over time. FedLU achieves this by combining two techniques: retroactive interference and passive decay. Retroactive interference allows the model to selectively unlearn specific knowledge embeddings related to outdated or irrelevant information, while passive decay gradually reduces the influence of such knowledge over time. This approach ensures that the KG remains up-to-date and relevant, adapting to changes in the underlying data distribution and facilitating the extraction of accurate and meaningful insights from the evolving knowledge graph.
@@
In the recommendation setting, RRL [38] is the initial work to apply reverse learning in recommendation. Based on the original recommendation objective, they devise a reverse objective. They train this newly proposed reverse objective to remove the impact of the target data. Selective and Collaborative Influence Function (SCIF) [39] is applied in recommendation to perform reverse gradient operations on the learned model. It directly removes the influence of target data by reverse operation, avoiding any kind of retraining. To further reduce the computational cost, it select the most pertinent user embedding for the reverse operation. Additionally, to enhance model utility, the collaborative data of the target data is considered as the collaboration preservation. This work quantifies the direct influence of unusable data on the optimization loss. However, the unusable data also affect the computational graph of recommender models such as LightGCN. This inspires a new Influence Function-based Recommendation Unlearning (IFRU) framework [45], focusing on the extension of the influence function to quantify the influence of unusable data in the computational graph aspect.
@@ 4.1 Unlearning Completeness
Membership Inference-based: One direction is to measure completeness based on attacking performance of membership inference (MI). Membership inference is a privacy attack that aims to determine whether a specific data point was part of the training dataset used to train a machine learning model. This becomes particularly relevant in machine unlearning and the success of MI can serve as a suitable measure for the quality of unlearning [21, 46]. To evaluate, one can conduct the membership inference for the removed data [26, 21, 18]. Ideally, the inference should show a high score for the original model, indicating the existence of these data in the training dataset. On the other hand, a low value should be presented for the unlearned model showing that these data is not used for training.
@@
Parameter Difference-based: A complete unlearning algorithm should have similar parameter weights with the retrained model. The difference can be measured as $d(\mathbf{W}^{u},\mathbf{W}^{r})$. A complete unlearning method should have a small distance measurement. Based on this assumption, [29] measures the Euclidean distance between the parameters returned by the unlearning algorithms and the parameters obtained via retraining from scratch. [34] measures the difference between normalized weight parameters.
Others: Several additional evaluation methods have been proposed. [29] intentionally modifies the target node’s label to an extra category and investigates whether the unlearned model classifies the node to this extra label. An extra binary feature is injected into the node features to help the model associate the deleted nodes with the extra-label category. Intuitively, a complete unlearning method should never predict the extra label. Similarly, [34] append an extra binary feature to all nodes and set the extra binary feature as 1/0 for the deleted nodes/other nodes respectively. They also add an extra category and change all deleted nodes to this extra category. They compare the weight norm of the injected channel before and after unlearning to determine whether the unlearning is complete. [33] propose an evaluation method based on adversarial attack. They add a certain number of edges to the training graph where each added edge links two nodes from different classes. The adversarial noise is expected to reduce the performance due to misleading the representation learning. Thereafter, they use the adversarial edges as unlearning requests and evaluate the utility of the estimated model. A larger utility gain indicated a higher unlearning efficacy. [35] uses the utility performance on the deleted data as a measurement for unlearning completeness. They assume that a smaller utility performance on the forgetting sets indicates that the model forgets more thoroughly.
Despite these commonly used measurements, there remains ongoing debate about how to best evaluate unlearning completeness. Some researchers argue that unlearning is only well-defined at the algorithmic level, where an entity’s only possible auditable claim to unlearning is that they used a particular algorithm designed to allow for external scrutiny during an audit [47].
@@ 4.3 Model Utility
Graph unlearning is inherently a multi-objective task. In addition to erasing the target data, the unlearning method needs to ensure that the performance on the remaining data is not compromised. In an extreme case, one can unlearn by randomly initializing the model but this will degrade the model performance significantly [29]. Therefore, preserving the utility performance of downstream graph tasks is also a crucial objective.
Model utility refers to the assessment of how effective and reliable an unlearning model is in achieving its intended purpose. To evaluate this aspect, common utility metrics such as F1 score [17], accuracy [31] for classification task, AUROC for link prediction [21], Normalized Discounted Cumulative Gain (NDCG) and Hit Ratio (HR) for recommendation task [48] are often employed. The specific utility metrics used depend on the nature of the downstream task. Model utility is frequently compared against the ‘scratch method’ (i.e., training from scratch), which, although less efficient, provides a baseline for utility performance evaluation.
@@ 5 Graph Unlearning Applications
Graph unlearning stands as an important topic in the realm of machine learning and data privacy. Due to the exponential development of graph-structured data across multiple domains, ensuring the right to be forgotten and protecting individual privacy have become of the utmost importance. As we observe the remarkable success of graph machine learning applications, it is important that we devise effective graph unlearning strategies. By enabling the removal of all forgotten data while preserving the integrity of interconnected networks, graph unlearning gives individuals greater control over their data and nurtures confidence in the responsible management of data. In the forthcoming sections, we first present potential unlearning requests from a user’s perspective in practice and then discuss the applications of graph unlearning in a variety of domains.
5.1 Real-world Graph Unlearning Examples from an Individual Perspective
Unlearning requests are increasingly common in various practical scenarios. Figure 3 illustrates several examples of unlearning requests for a single individual across different contexts. In one instance, the individual used an online dating application to seek a romantic relationship. After several dates, they found a significant other and subsequently decided to delete their account on the dating platform. The individual and their partner began watching movies together, enjoying the experience. However, following a breakup, the individual became dissatisfied with the platform’s recommendations of movies that are similar to ones that they had previously watched together, prompting an unlearning request to remove those viewing records. Another example, shown on the right-hand side, pertains to the healthcare domain. In this scenario, the individual needed to transfer hospitals due to relocation, resulting in the decision to move their health records to the new facility. In both scenarios, unlearning requests emerge from the user side and the models trained on the original data needs to unlearn the specific information in order to align with their requests for data removal or modification.
@@ 5.2 Social Networks
The application of graph unlearning in social networks is essential for mitigating the risk that a trained model may inadvertently learn sensitive information about users’ social relationships. On the basis of the underlying graph structure, machine learning models are frequently employed in social network analysis to infer user behavior, preferences, and social interactions. However, these models may inadvertently capture and encode sensitive information within their learned representations, such as private friendships, personal interests, and social circles [49]. Graph unlearning has the potential to provide means of mitigating this privacy concern. By employing graph unlearning techniques to the trained model, it is possible to delete or anonymize sensitive user data from the model’s learned parameters. This promotes a more secure and ethical environment for social network analysis by ensuring that the model’s predictions and insights are obtained without compromising the privacy and confidentiality of individual users. Through graph unlearning, researchers and organizations can strike a delicate equilibrium between leveraging the power of machine learning in understanding social dynamics and respecting and preserving users’ privacy rights in the ever-changing social network landscape [50, 51].
@@ 5.3 Confidentiality in Financial Networks
In financial transactions and fraud detection, learning models are frequently used to analyze transaction patterns and identify potential anomalies. However, these models may inadvertently acquire sensitive information, such as the financial histories, transaction details, and account balances of individuals, thereby jeopardizing the confidentiality and privacy of users’ financial data. By applying graph unlearning techniques to these models, it is possible to remove or obfuscate sensitive data from their learned representations. This ensures that the models can continue to provide accurate insights and detect fraudulent activities without compromising the privacy/confidentiality of individual users financial information [52]. Adopting graph unlearning in financial networks establishes a robust and ethical framework for data privacy and security, fostering user confidence and protecting sensitive financial data in this essential domain.
Figure 3: Real-world graph unlearning applications.

### v4

@@ 3.2.1 Convex Setting
A common assumption for theoretical guarantees of unlearning method is (strong) convexity, where the Hessian of the loss function with respect to the input is positive (semi-)definite. Rigorously, for loss function $L$, model weights $W_{1},W_{2}$, and a dataset $\mathcal{D}$, convexity is defined as:
$L(\lambda W_{1}+(1-\lambda)W_{2},\mathcal{D})\leq L(\lambda W_{1},\mathcal{D})+L((1-\lambda)W_{2},\mathcal{D}),$
for all $0\leq\lambda\leq 1$. For strong convexity, the weak inequality above becomes a strict less-than (<<). When the convexity assumption holds, the loss landscape of the model contains exactly one unique minimum, which is also a global minimum—making proving mathematical bounds vastly simpler. However, for this powerful condition, we trade off applicability: convexity condition is very strict, and mostly only apply on linear model with simple loss functions.
The paper [52] introduces a graph embedding technique called Graph Scattering Transform (GST). GST relies on a collection of multiresolution graph wavelets, a pointwise nonlinear activation function, and a low-pass operator. Building upon GST, [14] propose an interesting graph unlearning approach, aptly named Unlearning GST. By utilizing GST embeddings, Unlearning GST defines an unlearning model denoted as $\mathcal{M}^{\prime}$, which updates the trained model from $W$ to $W^{\prime}$. The latter represents an approximation of the unique optimizer of $L(W,\mathcal{D}^{\prime})$. The derivation of $W^{\prime}$ is as follows:
To begin, $H_{W}$ is defined as the Hessian of $L(.,\mathcal{D}^{\prime})$ at $W$ and is computed as:
@@
Using the obtained gradient difference, $W^{\prime}$ is derived as:
@@
$W^{\prime}$ thus represents the updated model obtained through this unlearning process. However, this approach has a limitation related to the computation of the wavelet transform, which is cubic for node removal and quadratic for node feature removal.
@@
The authors in [37] introduce another unlearning method, Graph Influence Function (GIF). GIF aims to model the influence of each training data point on the model with respect to various performance criteria and subsequently eliminates the negative impact. The method incorporates an additional loss term that considers the influence of neighbors and estimates parameter changes in response to a $\epsilon$-mass perturbation in deleted data. Note that while GIF can practically be applicable to the non-convex setting, the authors highlighted the decrease in its performance on sufficiently deep models as the setting deviates far enough from the theoretical assumptions. Another similar approach is “Projector” by Cong et al. [38] This method aims to remove the impact of deleted data from the learned model by projecting the weight parameters to a different subspace that is unrelated to the deleted data. Specifically, Projector is designed to work with linear GNNs, and demonstrates that all gradients lie within the linear span of all node features. Leveraging this insight, it performs unlearning using an orthogonal projection represented as a weighted combination of the remaining node features. By applying this technique, the learned model can effectively adapt to the removal request and update itself while minimizing the impact of the deleted data. IDEA [43] generalizes certified unlearning to all types of unlearning requests. Cognac [45] brings the contrastive approach to unlearning and applies them on nodes that are affected by the removal request, but requires the graph to be homophily and the model to be homophily-preserving.
@@ 3.2.2 Non-Convex Setting
In the pursuit of approximate unlearning within the non-convex setting, CEU is introduced in [36]. CEU specifically targets edge unlearning in graphs. To achieve this, CEU focuses on reversing the influence of the edge that needs to be removed from the new model $W^{\prime}$, leveraging the concept of influence functions [54]. The authors propose an estimation of the influence function by upweighting the set of all affected nodes and subsequently calculate the reverse of the Hessian matrix multiplied by the gradient vector to obtain the influence. To address computational complexity, they employ conjugate gradient optimization, effectively reducing the computational burden. ETR [46] follows a similar approach, but edit weights accordingly to the Fisher information matrix, which is computed approximately and efficiently.
@@
The unlearning process, as explored in [55], can potentially harm the performance of the underlying predictive model. In an effort to mitigate this issue and avoid updating the trained model, [16] introduces GNNDelete. This approach formalizes two essential properties for the GNN deletion method: deleted edge consistency and neighborhood influence. Deleted edge consistency ensures that the predicted probabilities for deleted edges in the unlearned model stay similar to those for nonexistent edges. Similarly, neighborhood influence guarantees that predictions in the local vicinity of the deletion retain their original performance and remain unaffected by the removal. In pursuit of efficiency and scalability, GNNDelete employs a layerwise deletion operator to modify a pretrained GNN model. Upon receiving deletion requests, GNNDelete freezes the existing model weights and introduces small, shared weight matrices across the nodes in the graph. GNNDelete guarantees strong performance by ensuring the difference between node representations obtained from the trained model $\mathcal{M}$ and those revised by GNNDelete $\mathcal{M}^{\prime}$ remains theoretically bounded.
Some approaches utilize knowledge distillation (KD) to unlearn specific subsets of the training data. Distill-to-Delete (D2DGN) [42] takes inspiration from KD and distill the full model to a new unlearned one, using loss terms to either destroy or retain knowledge of some training data. Specifically, D2DGN uses the cross-entropy, then maximizes it for the unlearning data and minimizes for the rest. GUKD [40] flips the traditional KD settings and instead applies distillation to the unlearning model with soft label from an independently retrained model on the remaining data. This retrained teacher model belongs to a smaller architecture to reduce computational overhead.
Instead of finding the most affected weights, which is the case for influence function-based methods, some work aim for detecting high influence nodes instead. MEGU [41] supports all type of removal requests by addressing the delete requests’ effects on these important nodes. This is done through simultaneously optimizing two loss terms, where one unlearns the requested-for-deletion nodes, and another preserves original utility of the model, on the important nodes only. AGU [49] separately addresses edge removal and feature removal, and in turn supports multiple types of graph unlearning by some combination of the two above. SGU [50] further improves on unlearning capability and performance by proposing a better node selection method named Node Influence Maximization, which is much faster than other competing baselines.
Newer works focus more on substantially reducing runtime for graph unlearning by simply not doing computation. Unlink-to-Unlearn (UtU) [44] prevents message passing of the unlearning nodes by removing unlearned edges during inference. UtU is very fast, and the authors conducted a thorough experiment to empirically verify its unlearning power. A more recent and notable method is ScaleGUN [50], which works on all graph unlearning request types and deal with the certified unlearning setting. It accomplishes this by determining loss bounds that would break the unlearning certification, and only commits lazy local propagation until when that bound is crossed.
@@
Regarding specialized approximate unlearning for specific applications, in the knowledge graph setting, KGU Schema [48] queries subgraphs similar to the removal request based on the graph entity schema, and only commit unlearning on them. For the recommendation setting, RRL [32] is the initial work to apply reverse learning in recommendation. Based on the original recommendation objective, they devise a reverse objective. They train this newly proposed reverse objective to remove the impact of the target data. Selective and Collaborative Influence Function (SCIF) [31] is applied in recommendation to perform reverse gradient operations on the learned model. It directly removes the influence of target data by reverse operation, avoiding any kind of retraining. To further reduce the computational cost, it select the most pertinent user embedding for the reverse operation. Additionally, to enhance model utility, the collaborative data of the target data is considered as the collaboration preservation. This work quantifies the direct influence of unusable data on the optimization loss. However, the unusable data also affect the computational graph of recommender models such as LightGCN. This inspires a new Influence Function-based Recommendation Unlearning (IFRU) framework [33], focusing on the extension of the influence function to quantify the influence of unusable data in the computational graph aspect.
@@ 4.1 Unlearning Completeness
Membership Inference-based: One direction is to measure completeness based on the performance of a membership inference attack (MIA). Membership inference is an attack that aims to determine whether a specific data point was part of the training dataset used to train a machine learning model. This becomes particularly relevant in machine unlearning and the success of MIA can serve as a suitable measure for the quality of unlearning [16, 56]. To evaluate, one can conduct the membership inference for the removed data. Ideally, the inference should show a high score for the original model, indicating the existence of these data in the training dataset. On the other hand, a low value should be presented for the unlearned model showing that these data is not used for training.
@@
Parameter Difference-based: A complete unlearning algorithm should have similar parameter weights with the retrained model. The difference can be measured as $d(\mathbf{W}^{u},\mathbf{W}^{r})$. Based on this assumption, [26] measures the Euclidean distance between the parameters returned by the unlearning algorithms and the parameters obtained via retraining from scratch. [38] measures the difference between normalized weight parameters.
Others: Several other less commonly used evaluation methods have been proposed. [26] intentionally modifies the target node’s label to an extra category and investigates whether the unlearned model classifies the node to this extra label. An extra binary feature is injected into the node features to help the model associate the deleted nodes with the extra-label category. Intuitively, a complete unlearning method should never predict the extra label. Similarly, [38] append an extra binary feature to all nodes and set the extra binary feature as 1/0 for the deleted nodes/other nodes respectively. They also add an extra category and change all deleted nodes to this extra category. They compare the weight norm of the injected channel before and after unlearning to determine whether the unlearning is complete. [37] propose an evaluation method based on adversarial attack. They add a certain number of edges to the training graph where each added edge links two nodes from different classes. The adversarial noise is expected to reduce the performance due to misleading the representation learning. Thereafter, they use the adversarial edges as unlearning requests and evaluate the utility of the estimated model. A larger utility gain indicated a higher unlearning efficacy. [58] uses the utility performance on the deleted data as a measurement for unlearning completeness, assuming a smaller utility performance on the forgetting sets indicates that the model forgets more thoroughly.
There remains ongoing debate about how to best evaluate unlearning completeness, arguing against the commonly-used measurements. For instance, MIA-based methods’ reliability depend on the quality of the underlying attack, which often fails for well-generalized models [59, 60], providing a false sense of security. Moreover, as MIAs are logit-based, a low attack success rate may not imply that the underlying latent space is unlearned, and only the shallow classification layers are modified instead [61]. Prediction Difference and Parameter Difference methods both have a very simple yet practical failure mode, that is when there are duplicates or dependency in the data [62]. Parameter Difference may prove to be unreliable, as functionally-equivalent models can have very different weights [63]. Some argue that unlearning is only well-defined at the algorithmic level, where the only possible auditable claim to unlearning is that they used a particular algorithm designed to allow for external scrutiny during an audit [64].
@@ 4.2 Unlearning Efficiency
Overall, most methods’ complexity scales linearly with the number of layers, and quadratic with the feature dimension [22].
@@ 4.3 Model Utility
Graph unlearning is inherently a multi-objective task. In addition to erasing the target data, the unlearning method needs to prevent performance degradation on the remaining data [26]. Therefore, preserving model utility for downstream graph tasks is also a crucial objective. Model utility refers to the assessment of how effective and reliable an unlearning model is in achieving its intended purpose. To evaluate this aspect, common utility metrics such as F1 score [15], accuracy [57] for classification task, AUROC for link prediction [16], Normalized Discounted Cumulative Gain (NDCG) and Hit Ratio (HR) for recommendation task [65] are often employed. The specific utility metrics used depend on the nature of the downstream task. Model utility is frequently compared against the ‘scratch method’ (i.e., training from scratch), which, although less efficient, provides a baseline for utility performance evaluation.
@@ 5 Graph Unlearning Applications
Graph unlearning stands as an important topic in the realm of machine learning and data privacy. Due to the exponential development of graph-structured data across multiple domains, ensuring the right to be forgotten and protecting individual privacy have become of the utmost importance. As we observe the remarkable success of graph machine learning applications, it is important that we devise effective graph unlearning strategies. By enabling the removal of all forgotten data while preserving the integrity of interconnected networks, graph unlearning gives individuals greater control over their data and nurtures confidence in the responsible management of data. In the next sections, we first present potential unlearning requests from a user’s perspective in practice and then discuss the applications of graph unlearning in a variety of domains.
5.1 Realworld Graph Unlearning from an Individual Perspective
Figure 2: Real-world graph unlearning applications.
Unlearning requests are increasingly common in practice. Figure 2 shows three examples for a single individual. In the first, the individual uses an online dating application to seek a relationship, matches with a partner, and then deletes their account. In the second, after the breakup, they request removal of movie records—primarily chosen by the partner—that they had watched together, so the platform stops recommending similar titles. In the third, the individual transfers to a new hospital and asks that their previous health records be removed from the original facility’s models. In each case, the request originates from the user, and the trained model must unlearn the specified information to comply.
@@ 5.2 Social Networks
The application of graph unlearning in social networks is essential for mitigating the risk that a trained model may inadvertently learn sensitive information about users’ social relationships. On the basis of the underlying graph structure, machine learning models are frequently employed in social network analysis to infer user behavior, preferences, and social interactions. However, these models may inadvertently capture and encode sensitive information within their learned representations, such as private friendships, personal interests, and social circles [66]. This is also the case for the e-commerce domain, where social networks can be used as input to a recommender system [67]. Graph unlearning has the potential to provide means of mitigating this privacy concern. By employing graph unlearning techniques to the trained model, it is possible to delete or anonymize sensitive user data from the model’s learned parameters. This promotes a more secure and ethical environment for social network analysis by ensuring that the model’s predictions and insights are obtained without compromising the privacy and confidentiality of individual users. Through graph unlearning, researchers and organizations can strike a delicate equilibrium between leveraging the power of machine learning in understanding social dynamics and respecting and preserving users’ privacy rights in the ever-changing social network landscape [68, 69].
@@ 5.3 Confidentiality in Financial Networks
In financial transactions and fraud detection, learning models are used to analyze transaction patterns and identify anomalies. However, they may also capture sensitive information such as financial histories, transaction details, and account balances, jeopardizing user privacy. Graph unlearning can remove or obfuscate this data from learned representations, enabling accurate insights and fraud detection without retaining sensitive financial information [70]. Adopting unlearning in financial networks provides a robust, ethical framework for privacy and security, fostering user confidence.